OpenStudy (anonymous):

4/x-2=-3/x+5 + 7/(x+5)(x-2) How do you go about solving this equation? Please show step by step. Thank you.

OpenStudy (anonymous):

both time (x+5)(x-2) in two sides, then we can have 4(x+5)=-3(x-2)+7 move -3(x-2) to right 4(x+5)+3(x-2)=7 spread it 4x+20+3x-6=7 keep only x one side 4x+3x=7-20+6 combin it 7x=-7 then x=-1

OpenStudy (anonymous):

move -3(x-2) to left

OpenStudy (anonymous):

so if you multiply the right side by x+5 does it take care of both the x+5's? if that makes sense?

OpenStudy (anonymous):

I'm trying to figure out what you have to do to get rid of the x+5 under the 7?

OpenStudy (anonymous):

multiply (x+5)(x-2) (x+5)(x-2) is one part

OpenStudy (anonymous):

not only (x+5) or (x-2)

OpenStudy (anonymous):

Okay I see I think ...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:)